# Parameter-Dependent Control Lyapunov Functions for Stabilizing Nonlinear Parameter-Varying Systems

### v1

@@ VI Simulation Results
We now test the proposed approach on a simple inverted pendulum system and a rocket landing problem. We used Matlab 2023a with YALMIP [25] and Mosek [26] to solve all the optimization problems.
VI-A Inverted pendulum with varying control effectiveness
We first consider a toy example: an inverted pendulum with states $x=[\phi,\dot{\phi}]^{T}$ and control input $u=\tau$. The dynamics are given by
$\dot{x}(t)=\begin{bmatrix}\dot{\phi}\\ \frac{g}{L}\sin(\phi)-\frac{b}{mL^{2}}\dot{\phi}\end{bmatrix}+\theta(t)\begin{bmatrix}0\\ \frac{1}{mL^{2}}\end{bmatrix}u(t),$ (36)
where $m=0.1$, $L=0.5$, $b=0.2$, and $g=9.81$ denote the mass, length, damping ratio, and gravitational constant, and $\left\lvert u(t)\right\rvert\leq 1$ for all $t\geq 0$. The time-varying parameter $\theta(t)$ is introduced to characterize varying control effectiveness due to actuator failure or nonlinearity. It is assumed that $\theta(t)\in\Theta=[0.2,1]$ and $|\dot{\theta}(t)|\leq 0.1$. Note that $\theta=0.2$ indicates a loss of 80% control effectiveness compared to the scenario with $\theta=1$.
@@ VI-A Inverted pendulum with varying control effectiveness
To find a PD-CLF, we approximated $\sin(\phi)$ with a third-order polynomial $\phi-\phi^{3}/6$, represented the dynamics in the linear-like form 18 and then solved the SOS optimization problem 32 to maximize the region of stabilization. We selected the decision variables as $X(\phi,\theta)$ and $Y(X,\theta)$, both of which has the order of 2, and the set $\mathcal{X}$ to be $[-2,2]\times[-2,2]$ The resulting PD-CLF is given by $V(x,\theta)=x^{T}X^{-1}(\phi,\theta)x$, where
$\small\begin{split}X(\phi,\theta)=&\begin{bmatrix}0.236&-0.405\\ -0.405&4.0\end{bmatrix}+\begin{bmatrix}-0.336&0\\ 0&0\end{bmatrix}\theta+\begin{bmatrix}0.84&-1.08\!\times\!10^{-5}\\ -1.08\!\times\!10^{-5}&0\end{bmatrix}\theta^{2}+\begin{bmatrix}1.9\!\times\!10^{-4}&0\\ 0&0\end{bmatrix}\phi^{2}\end{split}$
@@
after ignoring terms with very small coefficients. Note that solving the optimization problem 32 also gives a control law
$u(x,\theta)=Y(x,\theta)X^{-1}(\phi,\theta)x.$ (37)
@@
The PD region of stabilization (ROS) determined by $V(x,\theta)$ as $\Omega(V,x,\theta)=\{(x,\theta):V(x,\theta)\leq 1,\theta\in\Theta\}$ is illustrated in Fig. 2. Fig. 2 also includes the projection of $\Omega(V,x,\theta)$ onto the $x$-plane with different $\theta$ values as well as the robust ROS, which is obtained by removing the dependence of the decision variables $X$ and $Y$ on $\theta$ in solving the SOS problem 32. We can see that the robust ROS overlaps with the projection of the PD=ROS onto $\theta=0.2$, which is the smallest among all projected regions. By using the PD-ROS, we are able to enlarge the set of states from which the system can be stabilized by considering the initial values of $\theta$, i.e., $\theta(0)$.
@@
Figure 2: PD region of stabilization (PD-ROS) (left) and its projection onto $x$-planes with different $\theta$-values together with the robust ROS (right). Robust ROS overlaps with the smallest projection of the PD-ROS, corresponding to $\theta=0.2$
@@
Figure 3: Trajectories of states (left), Lyapunov function (top right) and control input (bottom right) starting from four different initial points under $\theta(t)=0.6+0.4\cos(0.2t)$ and the control law 37.
Figure 4: Trajectories of states (left), Lyapunov function (top right) and control input (bottom right) starting from four different initial points under $\theta(t)=0.6+0.4\cos(0.2t)$ and the min-norm control law from solving the QP problem.
VI-B Rocket landing in the presence of varying mass
@@
$\begin{bmatrix}\dot{y}\\ \dot{z}\\ \dot{\phi}\\ \dot{v}_{y}\\ \dot{v}_{z}\\ {\ddot{\phi}}\end{bmatrix}={\begin{bmatrix}v_{y}\\ v_{z}\\ \dot{\phi}\\ 0\\ g\\ 0\end{bmatrix}}+\begin{bmatrix}0&0\\ 0&0\\ 0&0\\ \frac{1}{m}&0\\ 0&\frac{1}{m/\theta}\\ -\frac{l\cos{\phi}}{J/\theta}&\frac{l\sin{\phi}}{J/\theta}\end{bmatrix}\begin{bmatrix}F_{y}\\ F_{z}\end{bmatrix},$ (38)
@@
where $y$, $z$ and $\phi$ are the position in $y$ axis, position in $z$ axis and the angle between the rocket body and upwards direction, $m=1$, $l=1$, and $J=\frac{1}{12}m(2l)^{2}$ denotes the mass, the length from center of mass (COM) to the thruster, and inertia of the rocket. Here we assume that the COM is in the middle of the rocket and is time-invariant to simplify the modeling. $g=9.81$ is the gravitational constant, and $F_{y}$ and $F_{z}$ denotes the thrust forces in the $y$ and $z$ axes, respectively. The parameter $\theta$ reflects the mass and inertia change due to fuel consumption, and is set to be $\theta(t)\in\Theta=[1,5]$, $\dot{\theta}(t)\in[0,0.1]$ for all $t\geq 0$. Note that $\theta=5$ corresponds to the mass and inertia reduced to 20% of those associated with $\theta=1$.
@@
Due to the constant $g$ in $f(x)$, the dynamics in 38 cannot be written in the linear-like polynomial form 18. To solve this issue, we define $F_{z}=\hat{F}_{z}+\frac{mg}{\theta}$, and rewrite the dynamics in the form of new control input $\hat{u}=[F_{y},\hat{F}_{z}]^{T}$, which can further be rewritten in the linear-like form 18. As a result, the synthesis procedure based on the SOS optimization problem 32 can be applied. In the procedure, we selected the Lyapunov and controller matrix variables as $X(\phi,\theta)$ and $Y(\phi,\theta)$, both of which have the order of 2. We decided not to include more dependent states in $X$ and $Y$ to limit the computational load. The local synthesis set $\mathcal{X}$ is selected to be $\mathcal{X}=\left\{x\in\mathbb{R}^{6}:y^{2}+z^{2}\leq 36,~{}\phi^{2}+(\dot{\phi})^{2}\leq(\frac{\pi}{3})^{2},~{}v_{y}^{2}+v_{z}^{2}\leq 4\right\}$. The control input constraint were set to be $\left\lvert F_{y}\right\rvert\leq 50$ and $\left\lvert\hat{F_{z}}\right\rvert\leq 50$.
@@
After getting the PD-CLF and the PD controller from solving 32, we simulated the system starting from initial states $x(0)=[y_{0},z_{0},\phi_{0},0,-1,0]^{T}$ under the trajectory $\theta(t)=1+0.08t$ the PD controller from solving 32. The results for three cases of $(y_{0},z_{0},\phi_{0})$, are shown in Fig. 5. Notice that $x(0)$ for all three cases are actually out of the set $\mathcal{X}$ used in the synthesis procedure. However, the controller was able to land the rocket successfully in all these cases.
Figure 5: Trajectories of states (left), Lyapunov function (middle two) and control inputs (bottom) starting from three different initial points under $\theta(t)=1+0.08t$ and the PD controller. The rocket body is illustrated by the blue rod, and the goal point (i.e., the origin) is denoted by $\star$.
VII Conclusions
@@
This paper presented an approach to gain-scheduled stabilization of nonlinear parameter-varying (NPV) systems through parameter-dependent (PD) control Lyapunov functions (CLFs). For control-affine polynomial NPV systems, it introduces convex conditions based on sum of squares (SOS) programming to jointly synthesize a PD-CLF and a stabilizing controller with and without input constraints. Simulation results validated the effectiveness of the proposed approach, demonstrating its potential for enhancing the stabilization of NPV systems. The proposed PD-CLF provides a way to control nonlinear systems with large operating envelopes. Compared to the existing LPV approach, it does not need linearization or over-approximation that is typically required to obtain an LPV system.
@@
Future work includes the relaxation of the assumption imposed in the synthesis procedure that input constraint should be symmetric, and consideration of potential errors in measuring and estimating the scheduling parameters that have been studied in the LPV setting [27, 28].

### v2

@@ V Simulation Results
We used Matlab 2023a with YALMIP [26] and Mosek [27] to solve all the optimization problems.
V-A Numerical example
We first consider a two-dimensional toy example
$\dot{x}=\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}\theta(t)x_{2}\\ 19.62\left(x_{1}-{x_{1}^{3}}/{6}\right)-8x_{2}\end{bmatrix}+\theta(t)\begin{bmatrix}0\\ 40\end{bmatrix}u.$ (37)
@@ V-A Numerical example
where $\theta(t)\in\Theta=[0.05,1]$ and $|\dot{\theta}(t)|\leq 0.1$. Note that $\theta=0.05$ indicates a loss of 95% control effectiveness compared to the scenario with $\theta=1$ from the input matrix. To find a PD-CLF, we represented the dynamics in the linear-like form 21 and then solved the SOS problem 35 to maximize the region of stabilization. We selected the decision variables as $X(x_{1},\theta)$ and $Y(x,\theta)$, both of which have the order of 2, and the set $\mathcal{X}$ to be $[-5,5]\times[-5,5]$. The resulting PD-CLF is given by $V(x,\theta)=x^{T}X^{-1}(x_{1},\theta)x$, where
$X(x_{1},\theta)=\begin{bmatrix}0.81&-1.79\\ -1.79&24.82\end{bmatrix}+\begin{bmatrix}-1.16&-3.32\\ -3.32&-0.34\end{bmatrix}\theta+10^{-4}\begin{bmatrix}7&-114\\ -114&172\end{bmatrix}x_{1}^{2}+\begin{bmatrix}12.12&-0.50\\ -0.50&0.18\end{bmatrix}\theta^{2}.$
@@
Note that solving the optimization problem 35 also gives a control law
$u(x,\theta)=Y(x,\theta)X^{-1}(x_{1},\theta)x.$ (38)
@@
The PD region of stabilization (ROS) determined by $V(x,\theta)$ as $\Omega(V,1)=\{(x,\theta):V(x,\theta)\leq 1,\theta\in\Theta\}$ is illustrated in Fig. 2. Fig. 2 also includes the projection of $\Omega(V,1)$ onto the $x$-plane with different $\theta$ values as well as the robust ROS, which is obtained by removing the dependence of the decision variables $X$ and $Y$ on $\theta$ in solving the SOS problem 35. We can see that the robust ROS almost overlaps with the projection of the PD-ROS onto $\theta=0.05$, which is the smallest among all projected regions. Using the PD-ROS leads to an enlarged set of states from which the system can be stabilized by considering the initial values of $\theta$, i.e., $\theta(0)$.
@@
Figure 2: PD region of stabilization (PD-ROS) (left) and its projection onto $x$-planes with different $\theta$-values together with the robust ROS (right). Robust ROS overlaps with the smallest projection of the PD-ROS, corresponding to $\theta=0.0.05$
@@
(a) Pre-defined control law 26
(b) Min-norm control law 8
Figure 3: Trajectories of states (top), Lyapunov function (middle) and control input (bottom) starting from four initial points under the pre-defined control law 26 (left) and the min-norm control law 8 (right).
V-B Rocket landing under varying mass and inertia
@@
$\begin{bmatrix}\dot{y}\\ \dot{z}\\ \dot{\phi}\\ \dot{v}_{y}\\ \dot{v}_{z}\\ {\ddot{\phi}}\end{bmatrix}={\begin{bmatrix}v_{y}\\ v_{z}\\ \dot{\phi}\\ 0\\ -g\\ 0\end{bmatrix}}+\begin{bmatrix}0&0\\ 0&0\\ 0&0\\ \frac{1}{m/\theta}&0\\ 0&\frac{1}{m/\theta}\\ -\frac{l\cos{\phi}}{J/\theta}&\frac{l\sin{\phi}}{J/\theta}\end{bmatrix}\begin{bmatrix}F_{y}\\ F_{z}\end{bmatrix},$ (39)
@@
where $y$, $z$ and $\phi$ are the position in $y$ axis, position in $z$ axis and the angle between the rocket body and upwards direction, $m=1$, $l=1$, and $J=\frac{m}{12}(2l)^{2}$ denotes the mass, the length from center of mass (COM) to the thruster, and inertia of the rocket. Here we assume that the COM is in the middle of the rocket and is time-invariant to simplify the modeling. $g=9.81$ is the gravitational constant, and $F_{y}$ and $F_{z}$ denotes the thrust forces in the $y$ and $z$ axes, respectively. The parameter $\theta$ reflects the mass and inertia change due to fuel consumption, and is set to be $\theta(t)\in\Theta=[1,5]$, $\dot{\theta}(t)\in[0,0.1]$ for all $t\geq 0$. Note that $\theta=5$ corresponds to the mass and inertia reduced to 20% of those associated with $\theta=1$.
(a) Pre-defined control law 26
(b) Min-norm control law 8
Figure 4: Trajectories of states (top) and control inputs (bottom) starting from four initial points under the pre-defined control law 26 (left) and the min-norm control law 8 (right).
@@
To remove the constant $g$ so that we could rewrite the dynamics in the linear-like form 21, we define $F_{z}=\hat{F}_{z}+\frac{mg}{\theta}$, and rewrite the dynamics in the form of new control input $\hat{u}=[F_{y},\hat{F}_{z}]^{T}$. With Taylor series expansion to approximate the trigonometric functions, the new dynamics can further be rewritten in the linear-like polynomial form 21. In solving the SOS problem 35, we selected the Lyapunov and controller matrix variables as $X(\phi,\theta)$ and $Y(\phi,\theta)$, both of which have the order of 2. We decided not to include more dependent states in $X$ and $Y$ to reduce the computational load. The local synthesis set $\mathcal{X}$ is selected to be $\mathcal{X}=\left\{x\in\mathbb{R}^{6}:y^{2}+z^{2}\leq 36,~{}\phi^{2}+(\dot{\phi})^{2}\leq(\frac{\pi}{3})^{2},~{}v_{y}^{2}+v_{z}^{2}\leq 4\right\}$.
@@
After getting the PD-CLF and the PD controller from solving 35, we simulated the system starting from initial states $x(0)=[y_{0},z_{0},\phi_{0},0,-0.5,0]^{T}$ under the trajectory $\theta(t)=1+0.08t$. The state and input trajectories for three cases of $(y_{0},z_{0},\phi_{0})$ under the predefined control law and min-norm control law, are shown in Fig. 4. Notice that $x(0)$ for all three cases is actually out of the set $\mathcal{X}$ used in the synthesis procedure. However, both control laws were able to land the rocket successfully in all these cases. Additionally, the magnitude of the min-norm control signal is significantly smaller compared to that of the predefined control law. This reduction offers notable advantages in certain scenarios, particularly in avoiding input saturation.
VI Conclusions
@@
This paper presented an approach to gain-scheduled stabilization of nonlinear parameter-varying (NPV) systems through parameter-dependent (PD) control Lyapunov functions (CLFs). For control-affine polynomial NPV systems, it introduces convex conditions based on sum of squares (SOS) programming to jointly synthesize a PD-CLF and a stabilizing controller. Simulation results validated the effectiveness of the proposed approach, demonstrating its potential for enhancing the stabilization of NPV systems. The proposed PD-CLF provides a way to control nonlinear systems with large operating envelopes. Compared to the existing LPV approach, it does not need linearization or over-approximation that is typically required to obtain an LPV system.
@@
Future work includes consideration of potential errors in measuring and estimating the scheduling parameters that have been studied in the LPV setting [28, 29].